OpenStudy (anonymous):

Is y=cos(x/2) a vertical shrink or a horizontal shrink?

OpenStudy (luigi0210):

Here are both graphs

OpenStudy (anonymous):

Thanks, I already know what they look like I just don't know when its a horizontal or vertical shrink

OpenStudy (jdoe0001):

the so-called "horizontal shift" or "transformations" for the trig functions, occur when their "regular period" gets interrupted, either expanded or shrunk so, sine and cosine have a "regular period" or \(2\pi\) and you can get it by $$ cos(x) = cos(1x)\\ period = \frac{2\pi}{1} $$

OpenStudy (jdoe0001):

now, what you have is really $$ cos\pmatrix{\frac{x}{2}} = cos\pmatrix{\frac{1}{2}x}\\ period = \frac{2\pi}{\frac{1}{2}} \implies 4\pi $$

OpenStudy (jdoe0001):

so, the "normally \(2\pi\)" period, gets affected by the coefficient in front of the "x" variable in the expression, in this case, what would otherwise be a "normal \(2\pi \)" got stretched to a \(4\pi\)

OpenStudy (anonymous):

I just wanted to know if it was a vertical or horizontal shrink I already know all that